OpenStudy (anonymous):

What formula do I use for this question? A newly married couple wants to purchase a new home. They find an account that earns 5.3% interest compounded monthly. How much do they need to invest now in order to have a $10,000 down payment in 18 months?

OpenStudy (anonymous):

is it.. A=P(1+rt) A=P(1+r/m)mt A=R((1/r/m)mt-1/(r/m))

ganeshie8 (ganeshie8):

\[\huge 10,000 = x(1+\frac{0.053}{12})^{18}\]

OpenStudy (anonymous):

I just need the formula to use. out of those three?

ganeshie8 (ganeshie8):

if \(x \) amount is invested now @ 5.3 % rate, after 18 months it becomes 10,000

ganeshie8 (ganeshie8):

ok lets see which formula fits

ganeshie8 (ganeshie8):

this works. A=P(1+r/m)mt replace r= .053, m = 12, mt=12

OpenStudy (anonymous):

thats the formula?

ganeshie8 (ganeshie8):

yep.. one typo.. mt=18. its not 12

OpenStudy (anonymous):

okay thanks:)

ganeshie8 (ganeshie8):

because, t = 1.5 here.. thus mt=18